Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Tue Feb 13 17:31:52 MET 2001) Minor failures
BBC8 is faultering (high levels, loss of locking and/or no_1pps). No other known problems.

Medicina (M.Bondi at Tue Feb 13 17:50:04 MET 2001) Success
No problem detected during the observation

Shanghai (Huang Xinyong at Wed Feb 14 00:27:47 MET 2001) Success
The observing was recorded with VLBA formater.

Effelsberg (A. Kraus at Wed Feb 14 11:05:10 MET 2001) Minor failures
The scans between UT0758-0808 were lost due to urgent technical work.

Urumqi (Liu Xiang at Wed Feb 14 11:12:26 MET 2001) Minor failures
breaks of the communication between FS computer and the antenna computer occured in very short time.

Noto (C.Trigilio at Thu Feb 15 18:49:26 MET 2001) Minor failures
Lost data from 10:39 to 11:17 for antenna in prelimit and delay of the operator at the tape change.

Westerbork (Tony Foley at Fri Feb 16 16:29:19 MET 2001) Success
no known problems

Jodrell_Bank (Alastair Gunn at Wed Feb 21 16:05:08 MET 2001) Success
No known problems.

Cambridge (Alastair Gunn at Wed Feb 21 16:05:32 MET 2001) Minor failures
FS crash at about 12:10, restarted at about 12:12. Data should be good. No other known problems.